OpenStudy (anonymous):

Given the function g(x) = -4x + 5, find g(-4).

OpenStudy (anonymous):

Plug -4 in for x. Simple.

OpenStudy (anonymous):

g(-4)=21

jimthompson5910 (jim_thompson5910):

g(x) = -4x + 5 g(-4) = -4(-4) + 5 ... replace EVERY x with -4 g(-4) = 16 + 5 g(-4) = 21

OpenStudy (anonymous):

oohhhhh..... so it already gives me the value of x... i just need to find the answer

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

g(-4) is the function value (or y value) when x = - 4

OpenStudy (anonymous):

21